OpenStudy (anonymous):

Word Problem: Arlene wishes to invest $5000. If she invests part at 7% simple interest, part at 6%, and receives a total of $332 after one year, how much does she invest at each rate?

OpenStudy (amistre64):

seven@7% = (332 - 5000(.06))/(.07 -.06)

OpenStudy (amistre64):

id say 3200 @ 7% and the rest at 6%

OpenStudy (amistre64):

a + b = t; find an equation for all the money split into 2 parts ax + by = z; determine an equation that accounts for the percents and the total amount at the end now eliminate one of the parts a + b = t; *-x ax + by = z -ax -bx = -tx ax+by = z ------------- b(y-x) = (z -tx) b@y = (z-tx)/(y-x)